# Geometric Representation Condition Improves Equivariant Molecule Generation

### v1

@@ Appendix D Additional Experiment Results
Next, we investigate the impact of different pre-trained encoders, which could vary in model structure and proxy tasks used for pre-training. Specifically, we compare Unimol (Zhou et al., 2023), which employs a message-passing neural network framework incorporating distance features (i.e., DisGNN in (Li et al., 2023; 2024b)) and primarily uses naive coordinates denoising, with Frad (Feng et al., 2023), which adopts TorchMD (Thölke & De Fabritiis, 2022) as the backbone and utilizes carefully designed hybrid-denoising tasks. Both Unimol (Zhou et al., 2023) and Frad (Feng et al., 2023) are pre-trained on the GEOM-DRUG dataset until convergence. We visualize the T-SNE of the representations generated for GEOM-DRUG. As shown in Figure 6, the T-SNE of the Unimol representations exhibits a clearer clustering pattern based on node numbers compared to the Frad representations, which may suggest better representation learning. To further investigate, we utilize both encoders to train GeoRCG and subsequently evaluate the quality of molecule generation. The Frad-based GeoRCG achieves a Validity of $96.9(0.44)$ and Atom Stability of $84.4(0.27)$, while the Unimol-based GeoRCG achieves a Validity of $98.5(0.12)$ and Atom Stability of $84.3(0.12)$: Although the Frad-based GeoRCG produces slightly higher atom stability, its high variance and significantly lower validity suggest inferior performance. These findings, along with our main results, offer insights into the types of representations more effective for guiding molecule generation, suggesting that sensitivity to molecule size may be a critical factor.
@@
Quality of Conditionally Generated Molecules
@@ Appendix E Theoretical Analysis
In this section, we use SDE and score matching formulations of diffusion models to present our theoretical results, given their equivalence with the DDPM family (Song et al., 2021). We consider the random variable $x\in\mathbb{R}^{d}$, and use $q(\cdot)$ to denote the ground truth distributions, $p(\cdot)$ to denote the posterior distribution predicted by diffusion models. For instance, $q(x)$ is the ground truth distribution of the underlying data $x$, while $p_{\varphi}(r)$ is the predicted distribution of latent representations. We use $T$ to denote the total time of diffusion models, and $N$ to represent the discretization step number. We consider a SDE with continuous time $[0,T]$, as well as its discretized DDPM which has $N$ diffusion steps with step size $h:=T/N$. The forward process is denoted as $(x_{t})_{t\in[0,T]}\sim q_{t}$, and the reverse process is denoted as $(\bar{x}_{t})_{t\in[0,T]}\sim p_{t}$. If the reverse process is predicted by the score matching network, we use its parameters as the subscript. Please note that there are two different initialization of the reverse process: the end of forward process $q_{T}$ and standard Gaussian noise $\gamma^{d}$. We use superscript $q_{T}$ to differentiate the former from the latter.
@@ E.1.1 Provable Generation of Representations
Suppose Assumption E.1, Assumption E.2, Assumption E.3 hold, and the step size $h:=T/N$ satisfies $h\preceq 1/L_{r}$. Then the following holds,
@@
This is a direct conclusion from (Chen et al., 2023). In typical DDPM implementation, we choose $h=1$ and thus $T=N$. Remarkably, Proposition E.1 indicates the benefit of generating the representation first: since $d_{r}\ll d$, the generation quality (measured by the TV distance in Proposition E.1) of the low-dimensional representation can easily outperform directly generating the high-dimensional data points $x$. The theorem also accounts for applying a lightweight MLP as the denoising network while in the stage of generating the representation.
@@ E.1.2 Provable Second-Stage Generation
Suppose Assumption E.4, Assumption E.5, Assumption E.6 hold, and the step size $h:=T/N$ satisfies $h\preceq 1/L_{x}$. Suppose we sample $x\sim p_{\theta}(x|r)$ from Gaussian noise where $r\sim p_{\varphi}(r)$, and denote the final distribution of $x$ as $p_{\theta,\varphi}(x)$. Then the following holds,
${\rm TV}(p_{\theta,\varphi}(x),q(x))$ $\preceq\underbrace{\sqrt{{\rm KL}(q(x)||\gamma^{d+3})}\exp(-T)}_{\text{convergence of forward process}}+\underbrace{(L_{x}\sqrt{(d+3)h}+L_{x}m_{x}h)\sqrt{T}}_{\text{discretization error}}+\underbrace{\epsilon_{\theta,\rm score}\sqrt{T}}_{\text{score estimation error}}$ (23)
@@
${\rm TV}(p_{0},p_{0}^{q_{T}})\leq{\rm TV}(\gamma^{d+3},q_{T})\leq\sqrt{{\rm KL}(q(x)||\gamma^{d+3})}\exp(-T)$ (25)
@@
${\rm TV}(p_{0}^{q_{T}},q_{0})^{2}\leq{\rm KL}(q_{0}||p_{0}^{q_{T}})\preceq(\epsilon_{\theta,\rm score}^{2}+L_{x}^{2}(d+3)h+L_{x}^{2}m_{x}^{2}h^{2})T$ (26)
@@
$\big{|}\big{|}\nabla\ln\frac{q_{kh}}{q_{t}}(x_{kh})\big{|}\big{|}^{2}\preceq L_{x}^{2}(d+3)h+L_{x}^{2}h^{2}||x_{kh}||^{2}+(L_{x}^{2}+1)h^{2}||\nabla\ln q_{t}(x_{kh})||^{2}$ (32)
@@
$\preceq$ $\epsilon_{\theta,\rm score}^{2}+L_{x}^{2}(d+3)h+L_{x}^{2}h^{2}\mathbb{E}_{q_{0}}[||x_{kh}||^{2}]+L_{x}^{2}h^{2}\mathbb{E}_{q_{0}}[||\nabla\ln q_{t}(x_{t})||^{2}]+L_{x}^{2}\mathbb{E}_{q_{0}}[||x_{kh}-x_{t}||^{2}]$ (36)
$\preceq$ $\epsilon_{\theta,\rm score}^{2}+L_{x}^{2}(d+3)h+L_{x}^{2}h^{2}(d+3+m_{x}^{2})+L_{x}^{3}(d+3)h^{2}+L_{x}^{2}(m_{x}^{2}h^{2}+(d+3)h)$ (37)
$\preceq$ $\epsilon_{\theta,\rm score}^{2}+L_{x}^{2}(d+3)h+L_{x}^{2}h^{2}m_{x}^{2}$ (38)
Analogous to (Chen et al., 2023), using properties of Brownian motions and local martingales, we can apply Girsanov’s theorem and complete the stochastic integration. Since $q_{0}$ is the end of the reverse SDE, by the lower semicontinuity of the KL divergence and the data-processing inequality, we take the limit an+d obtain
${\rm KL}(q_{0}||p_{0}^{q_{T}})\preceq(\epsilon_{\theta,\rm score}^{2}+L_{x}^{2}(d+3)h+L_{x}^{2}h^{2}m_{x}^{2})T$ (39)
@@
${\rm TV}(p_{\theta,\varphi}(x),q(x))\preceq$ $\underbrace{\sqrt{{\rm KL}(q_{0|\varphi}||\gamma^{d+3})}\exp(-T)}_{\text{convergence of forward process}}+\underbrace{(L_{x}\sqrt{(d+3)h}+L_{x}m_{x}h)\sqrt{T}}_{\text{discretization error}}$ (41)
@@
To conclude this subsection, we provide a detailed characterization of the generalization error of representation-conditioned diffusion models. It is important to note that some parameters in our assumptions, such as Lipschitz scores and estimation errors, are not constants; they are functions of the SDE total time $T$ and the number of diffusion steps $N$. Our conclusions also explain why representation-conditioned generation methods remain competitive even when the number of second-stage diffusion steps $N$ is decreased for faster generation. This is because the score estimation error can remain small even when the number of diffusion steps is reduced, given the guidance from representations. As a result, reducing $N$ causes a slower increase in $\epsilon_{\varphi,\theta,\rm cond}(N)$ compared to the score estimation error without representation conditioning, leading to representation-conditioned generative models’ strong performance with fewer steps.

### v2

@@ Appendix D Additional Experiment Results
Correlation between molecule representations and node numbers
See Figure 4.
Additional Comparison with 2D&3D methods
We further compare GeoRCG with recent 2D&3D methods, including MiDi (Vignac et al., 2023) and LDM-3DG (You et al., 2023), which jointly learn and generate both 2D bond information and 3D geometry. However, these models rely on external tools like RDKit or OpenBabel (O’Boyle et al., 2011) for bond computation at the input stage, which enables them to implicitly leverage domain knowledge of such tools. In contrast, purely 3D methods like GeoRCG and those in our primary comparison utilize a coarse look-up table for bond estimation, which, while reflecting 3D learning, results in less accurate bond calculation, potentially biasing the comparison. Therefore, to provide a reference, we also report the performance of our models combined with the same external tools (e.g., OpenBabel) for precise bond calculations in our generated 3D conformations. Furthermore, as GeoRCG essentially captures 3D geometric distributions, we place more emphasis on 3D metrics that directly evaluate 3D learning capabilities, including BondLengthW1 and BondAngleW1 proposed by Vignac et al. (2023) and detailed in Appendix C. The results in Table 7 demonstrate that GeoRCG not only significantly outperforms MiDi and LDM-3DG on 3D metrics, highlighting the advantages of using a pure 3D model for learning 3D structures, but also further enhances EDM’s performance, which has already shown considerable promise in 3D learning.
Balancing Controllability
We proceed to investigate the “Balancing Controllablility” feature of GeoRCG introduced in Section 3.2. To this end, we conducted a grid search by varying both $w$ and $\mathcal{T}$ on QM9 dataset, as depicted in Figure 5. The results indicate a clear trend: increasing $w$ and decreasing $\mathcal{T}$ improve validity and stability at the expense of uniqueness, allowing for fine-grained, flexible control over molecule generation. At its best, this approach achieves a molecule stability of 93.9% and a validity of 97.42%, approaching the dataset’s upper bound, with a trade-off in lower uniqueness&validity of 86.82%.
Data $\sim$0.1 $\sim$0 98.7 99.8 98.9 99.9
LDM-3DG 3.56 0.2 94.03 99.38 94.89 97.03
EDM 0.44 0.1 90.7 99.2 91.7 98.5
EDM + OBabel 0.44 0.1 97.9 99.8 99.0 98.5
@@
Next, we investigate the impact of different pre-trained encoders, which could vary in model structure and proxy tasks used for pre-training. Specifically, we compare Unimol (Zhou et al., 2023), which employs a message-passing neural network framework incorporating distance features (i.e., DisGNN in (Li et al., 2023, 2024b)) and primarily uses naive coordinates denoising, with Frad (Feng et al., 2023), which adopts TorchMD (Thölke & De Fabritiis, 2022) as the backbone and utilizes carefully designed hybrid-denoising tasks. Both Unimol (Zhou et al., 2023) and Frad (Feng et al., 2023) are pre-trained on the GEOM-DRUG dataset until convergence. We visualize the t-SNE of the representations generated for GEOM-DRUG. As shown in Figure 6, the t-SNE of the Unimol representations exhibits a clearer clustering pattern based on node numbers compared to the Frad representations, which may suggest better representation learning. To further investigate, we utilize both encoders to train GeoRCG and subsequently evaluate the quality of molecule generation. The Frad-based GeoRCG achieves a Validity of $96.9(0.44)$ and Atom Stability of $84.4(0.27)$, while the Unimol-based GeoRCG achieves a Validity of $98.5(0.12)$ and Atom Stability of $84.3(0.12)$: Although the Frad-based GeoRCG produces slightly higher atom stability, its high variance and significantly lower validity suggest inferior performance. These findings, along with our main results, offer insights into the types of representations more effective for guiding molecule generation, suggesting that sensitivity to molecule size may be a critical factor.
@@
Ablation Study: Representation Perturbation
@@ Appendix E Theoretical Analysis
In this section, we use SDE and score matching formulations of diffusion models to present our theoretical results, given their equivalence with the DDPM family (Song et al., 2021). We consider the random variable $x\in\mathbb{R}^{N\times(d+3)}$, and use $q(\cdot)$ to denote the ground truth distributions, $p(\cdot)$ to denote the posterior distribution predicted by diffusion models. For instance, $q(x)$ is the ground truth distribution of the underlying data $x$, while $p_{\varphi}(r)$ is the predicted distribution of latent representations. We use $T$ to denote the total time of diffusion models, and $N_{d}$ to represent the discretization step number. We consider a SDE with continuous time $[0,T]$, as well as its discretized DDPM which has $N_{d}$ diffusion steps with step size $h:=T/N_{d}$. The forward process is denoted as $(x_{t})_{t\in[0,T]}\sim q_{t}$, and the reverse process is denoted as $(\bar{x}_{t})_{t\in[0,T]}\sim p_{t}$. If the reverse process is predicted by the score matching network, we use its parameters as the subscript. Please note that there are two different initialization of the reverse process: the end of forward process $q_{T}$ and standard Gaussian noise $\gamma^{d}$. We use superscript $q_{T}$ to differentiate the former from the latter.
@@ E.1.1 Provable Generation of Representations
Suppose Assumption E.1, Assumption E.2, Assumption E.3 hold, and the step size $h:=T/N_{d}$ satisfies $h\preceq 1/L_{r}$. Then the following holds,
@@
Here $a(\cdot)\preceq b(\cdot)$ means there exists a constant $C$ such that $a(\cdot)\leq Cb(\cdot)$ always holds. This is a direct conclusion from (Chen et al., 2023). In typical DDPM implementation, we choose $h=1$ and thus $T=N_{d}$. Remarkably, Proposition E.4 indicates the benefit of generating the representation first: since $d_{r}\ll d$, the generation quality (measured by the TV distance in Proposition E.4) of the low-dimensional representation can easily outperform directly generating the high-dimensional data points $x$. The theorem also accounts for applying a lightweight MLP as the denoising network while in the stage of generating the representation.
@@ E.1.2 Provable Second-Stage Generation
Suppose Assumption E.6, Assumption E.7, Assumption E.8 hold, and the step size $h:=T/N_{d}$ satisfies $h\preceq 1/L_{x}$. Suppose we sample $x\sim p_{\theta}(x|r)$ from Gaussian noise where $r\sim p_{\varphi}(r)$, and denote the final distribution of $x$ as $p_{\theta,\varphi}(x)$. Then the following holds,
${\rm TV}(p_{\theta,\varphi}(x),q(x))$ $\preceq\underbrace{\sqrt{{\rm KL}(q(x)||\gamma^{N(d+3)})}\exp(-T)}_{\text{convergence of forward process}}+\underbrace{(L_{x}\sqrt{N(d+3)h}+L_{x}m_{x}h)\sqrt{T}}_{\text{discretization error}}+\underbrace{\epsilon_{\theta,\rm score}\sqrt{T}}_{\text{score estimation error}}$ (25)
@@
${\rm TV}(p_{0},p_{0}^{q_{T}})\leq{\rm TV}(\gamma^{N(d+3)},q_{T})\leq\sqrt{{\rm KL}(q(x)||\gamma^{N(d+3)})}\exp(-T)$ (27)
@@
${\rm TV}(p_{0}^{q_{T}},q_{0})^{2}\leq{\rm KL}(q_{0}||p_{0}^{q_{T}})\preceq(\epsilon_{\theta,\rm score}^{2}+L_{x}^{2}N(d+3)h+L_{x}^{2}m_{x}^{2}h^{2})T$ (28)
@@
$\big{|}\big{|}\nabla\ln\frac{q_{kh}}{q_{t}}(x_{kh})\big{|}\big{|}^{2}\preceq L_{x}^{2}N(d+3)h+L_{x}^{2}h^{2}||x_{kh}||^{2}+(L_{x}^{2}+1)h^{2}||\nabla\ln q_{t}(x_{kh})||^{2}$ (34)
@@
$\preceq$ $\epsilon_{\theta,\rm score}^{2}+L_{x}^{2}N(d+3)h+L_{x}^{2}h^{2}\mathbb{E}_{q_{0}}[||x_{kh}||^{2}]+L_{x}^{2}h^{2}\mathbb{E}_{q_{0}}[||\nabla\ln q_{t}(x_{t})||^{2}]+L_{x}^{2}\mathbb{E}_{q_{0}}[||x_{kh}-x_{t}||^{2}]$ (38)
$\preceq$ $\epsilon_{\theta,\rm score}^{2}+L_{x}^{2}N(d+3)h+L_{x}^{2}h^{2}(N(d+3)+m_{x}^{2})+L_{x}^{3}N(d+3)h^{2}+L_{x}^{2}(m_{x}^{2}h^{2}+N(d+3)h)$ (39)
$\preceq$ $\epsilon_{\theta,\rm score}^{2}+L_{x}^{2}N(d+3)h+L_{x}^{2}h^{2}m_{x}^{2}$ (40)
Analogous to (Chen et al., 2023), using properties of Brownian motions and local martingales, we can apply Girsanov’s theorem and complete the stochastic integration. Since $q_{0}$ is the end of the reverse SDE, by the lower semicontinuity of the KL divergence and the data-processing inequality (Beaudry & Renner, 2011), we take the limit and obtain
${\rm KL}(q_{0}||p_{0}^{q_{T}})\preceq(\epsilon_{\theta,\rm score}^{2}+L_{x}^{2}N(d+3)h+L_{x}^{2}h^{2}m_{x}^{2})T$ (41)
@@
${\rm TV}(p_{\theta,\varphi}(x),q(x))\preceq$ $\underbrace{\sqrt{{\rm KL}(q_{0|\varphi}||\gamma^{N(d+3)})}\exp(-T)}_{\text{convergence of forward process}}+\underbrace{(L_{x}\sqrt{N(d+3)h}+L_{x}m_{x}h)\sqrt{T}}_{\text{discretization error}}$ (43)
@@
Extension to Equivariant Diffusion Models.
The previous conclusions are generic and can be applied to general representation-conditioned generation. However, so far we only consider traditional diffusion models without taking into account the permutation $\Pi$ and SE(3) transformation $\Omega$ invariance/equivariance of the diffusion model. We thus extend our theory specifically to equivariant diffusion models that operate on symmetry structures, which is the case for our experiments. Moreover, the previous results assume that both the diffusion process and the denoising model treat atom coordinates $\mathbf{x}$ and atom type features $\mathbf{h}$ identically, while the fact is that (1) the noises of atom coordinates always have zero center of mass (CoM), thus actually lies in a subspace with degree of freedom $3(N-1)$ as opposed to $3N$; (2) in the forward process, $\mathbf{x}_{t}$ and $\mathbf{h}_{t}$ are conditional independent give $\mathbf{x}_{0}$ and $\mathbf{h}_{0}$, and since the denoising network processes coordinates $\mathbf{x}_{t}$ and $\mathbf{h}_{t}$ differently, the score estimation error term can be further decomposed as in Assumption E.13.
Assumption E.13.
(fine-grained representation-conditioned score estimation error of second-stage diffusion). For all $t\in[0,T]$,
$\mathbb{E}_{r\sim p_{\varphi}(r),x_{t}\sim q_{t}(x_{t}|r)}[||s_{\theta}(x_{t},t,r)-\nabla\ln q_{t}(x_{t}|r)||^{2}]$ (46)
$=$ $\mathbb{E}_{r\sim p_{\varphi}(r),x_{t}\sim q_{t}(x_{t}|r)}[||s_{\theta}^{(\mathbf{x})}(x_{t},t,r)-\nabla_{\mathbf{x}_{t}}\ln q_{t}(\mathbf{x}_{t}|r)||^{2}+||s_{\theta}^{(\mathbf{h})}(x_{t},t,r)-\nabla_{\mathbf{h}_{t}}\ln q_{t}(\mathbf{h}_{t}|r)||^{2}]$ (47)
$\leq$ $(\epsilon_{\varphi,\theta,\rm cond}^{\mathbf{x}})^{2}+(\epsilon_{\varphi,\theta,\rm cond}^{\mathbf{h}})^{2}$ (48)
where $s_{\theta}^{(\mathbf{x})}(\cdot)$ and $s_{\theta}^{(\mathbf{h})}(\cdot)$ refer to the predicted score of coordinates and atom type features by the score network, respectively.
Combining all the pieces, we conclude the following result Theorem E.14 for equivariant diffusion models that generate 3D coordinates as described above.
Theorem E.14.
[…]
${\rm TV}(\tilde{p}_{\theta,\varphi}(x),\tilde{q}(x)):=$ $\alpha(p_{\theta,\varphi},\Pi,\Omega){\rm TV}(p_{\theta,\varphi}(x),q(x))$ (49)
$\preceq$ $\alpha(p_{\theta,\varphi},\Pi,\Omega)\Bigg{(}\underbrace{\sqrt{{\rm KL}(q_{0|\varphi}||\gamma^{N(d+3)}_{0-CoM})}\exp(-T)}_{\text{convergence of forward process}}+\underbrace{(L_{x}\sqrt{N(d+3)h}+L_{x}m_{x}h)\sqrt{T}}_{\text{discretization error}}$ (50)
$\ +\underbrace{(\epsilon_{\varphi,\theta,\rm cond}^{\mathbf{x}}+\epsilon_{\varphi,\theta,\rm cond}^{\mathbf{h}})\sqrt{T}}_{\text{conditional score estimation error}}+\underbrace{{\rm TV}(q_{0|\varphi},q_{0})}_{\text{representation generation error}}\Bigg{)}$ (51)
where $\alpha(\cdot)\in[0,1]$.
Proof.
The proof parallels the proof of Theorem E.12 with three distinct parts.
  1. 1.
    Due to the existence of the permutation group $\Pi$ and the SE(3) group $\Omega$, we need to consider the distribution $\tilde{p}_{\theta,\varphi}(x)$. Note the definition $\alpha(p_{\theta,\varphi},\Pi,\Omega):=\frac{{\rm TV}(\tilde{p}_{\theta,\varphi}(x),\tilde{q}(x))}{{\rm TV}(p_{\theta,\varphi}(x),q(x))}$, and by data processing inequality (Beaudry & Renner, 2011), $\alpha(p_{\theta,\varphi},\Pi,\Omega)\in[0,1]$. Specifically, when the denoising model $p_{\theta,\varphi}$ is constructed invariant/equivariant to permutations $\Pi$ and SE(3) transformations $\Omega$ (which means the model treats all elements in one equivalent class the same), $\alpha(p_{\theta,\varphi},\Pi,\Omega)$ reaches the minimum; see (You et al., 2023) for further explanations.
  2. 2.
    Since the Gaussian noises for coordinates are sampled from a subspace with zero center of mean, the prior distribution $\gamma^{N(d+3)}$ in Theorem E.12 should be replaced with $\gamma^{N(d+3)}_{0-CoM}$, the $N(d+3)$-dimensional Gaussian but with zero center of mass in the $N\times 3$-dimensional subspace for coordinates. It is notable that the degree of freedom of $\gamma^{N(d+3)}_{0-CoM}$ is actually $N(d+3)-3$, and the remaining of the proof still holds (Feng et al., 2024).
  3. 3.
    As explained before, in a properly designed forward process, $\mathbf{x}_{t}$ and $\mathbf{h}_{t}$ can be conditional independent give $\mathbf{x}_{0}$ and $\mathbf{h}_{0}$. Meanwhile, the denoising network processes coordinates $\mathbf{x}_{t}$ and $\mathbf{h}_{t}$ differently, the score estimation error term can be further decomposed as in Assumption E.13; see (Feng et al., 2024) for more details. Therefore, the term $\epsilon_{\varphi,\theta,\rm cond}\sqrt{T}$ in Theorem E.12 can be replaced by $\sqrt{(\epsilon_{\varphi,\theta,\rm cond}^{\mathbf{x}})^{2}+(\epsilon_{\varphi,\theta,\rm cond}^{\mathbf{h}})^{2}}$, which is further bounded by $(\epsilon_{\varphi,\theta,\rm cond}^{\mathbf{x}}+\epsilon_{\varphi,\theta,\rm cond}^{\mathbf{h}})$.
@@
In conclusion, we provide a detailed characterization of the generalization error of representation-conditioned diffusion models in this subsection. It is important to note that some parameters in our assumptions, such as Lipschitz scores and estimation errors, are not constants; they are functions of the SDE total time $T$ and the number of diffusion steps $N_{d}$. Our conclusions also explain why representation-conditioned generation methods remain competitive even when the number of second-stage diffusion steps $N_{d}$ is decreased for faster generation. This is because the score estimation error can remain small even when the number of diffusion steps is reduced, given the guidance from representations. As a result, reducing $N_{d}$ causes a slower increase in $\epsilon_{\varphi,\theta,\rm cond}(N_{d})$ compared to the score estimation error without representation conditioning, leading to representation-conditioned generative models’ strong performance with fewer steps.